OpenStudy (anonymous):

How could I solve this using substitution in systems of equations? 4x+5y=11 y=3x-13

OpenStudy (anonymous):

Ok so you sub in the: y = 3x - 13 into the first equation. 4x + 5(3x -13) = 11 4x + 15x - 65 = 11 19x -65 = 11 19x = 76 x = 4 and then you continue with the y-value :D